Solve the following differential equation by changing the independent variable:
x.d^2y/dx^2+(4x^2-1)dy/dx+4x^3=2x^3
1
Expert's answer
2020-06-08T19:41:55-0400
"x{d^2y\over dx^2}+(4x^2-1){dy\over dx}+4x^3y=2x^3""{d^2y\over dx^2}+(4x-{1\over x}){dy\over dx}+4x^2y=2x^2,x\not=0"

Here

"P=4x-\dfrac{1}{x},"

"Q=4x^2"

"R=2x^2"

Changing the independent variable from "x" to "z"


"{d^2y\over dz^2}+P_1{dy\over dz}+Q_1y=R_1"

where


"P_1=\dfrac{{d^2z\over dx^2}+P{dz\over dx}}{({dz\over dx})^2}"

"Q_1=\dfrac{Q}{({dz\over dx})^2}"

"R_1=\dfrac{R}{({dz\over dx})^2}"

We choose "z" such that


"Q_1=\dfrac{Q}{({dz\over dx})^2}=constant"

Let


"Q_1=\dfrac{4x^2}{({dz\over dx})^2}=1"

Then


"{dz\over dx}=2x"

Integrating

"z=x^2"

We have


"{d^2z\over dx^2}=2"

"P_1=\dfrac{{d^2z\over dx^2}+P{dz\over dx}}{({dz\over dx})^2}=\dfrac{2+(4x-\dfrac{1}{x})(2x)}{(2x)^2}=2"

"R_1=\dfrac{R}{({dz\over dx})^2}=\dfrac{2x^2}{(2x)^2}={1\over 2}"

"{d^2y\over dz^2}+2{dy\over dz}+y={1\over 2}"

Method of undetermined coefficients

Homogeneous second order differentional equation


"{d^2y\over dz^2}+2{dy\over dz}+y=0"

Characteristic equation


"\lambda^2+2\lambda+1=0""\lambda_1=\lambda_2=-1"

The general solution of the homogeneous differential equation is


"y_0=C_1ze^{-z}+C_2e^{-z}"

Let "Y=A," then


"{dY\over dz}=0, {d^2Y\over dz^2}=0"

"Y={1\over 2}"

The general solution of the nonhomogeneous equation


"y(z)=y_0+Y"

"y(z)=C_1ze^{-z}+C_2e^{-z}+{1\over 2}"

Substitute "z=x^2" and obtain the general solution


"y(z)=C_1x^2e^{-x^2}+C_2e^{-x^2}+{1\over 2}"